What is your average rate if you drive to campus averaging 50 mph and return home on the same route averaging 40 mpg
Answer by Alan3354(69443)

What is your average rate if you drive to campus averaging 50 mph and return home on the same route averaging 40 mpg
--------------
It's not 45, the avg of 50 & 40 because more time is spent going 40.
----
There's a formula similar to parallel resistance and parallel flows
Avg = 2*50*40/(50 + 40) = 4000/90 = 400/9
= 44.444... mi/hr (not mpg)
--------
I can derive the formula if you like.